If I have 43000 and I invest part of it on bonds paying 4% interest Annual and some to life inc paying 8% interest annual. And the total interest from this investment after 1year is 2880 how much did I invest in bonds
Answer by Boreal(15235)

x=invested at 4%
43000-x=invested at 8%
.04x+.08(43000-x)=2880
.04x+3440-.08x=2880
-.04x=-560
multiply both sides by -25
x=$14,000 in bonds@4%=$560
43000-x in life insurance @ 8% is 29,000@8%=$2320
They add to $2880.